OpenStudy (anonymous):

I need help to Solve 2x2 - 8x = -7.my choices are. negative 2 plus or minus square root of 2 negative 2 plus or minus 2 square root of 2 quantity of 2 plus or minus square root of 2 all over 2 2 plus or minus square root of 2 end root over 2

OpenStudy (anonymous):

By 2x2, do you mean \[2x ^{2}\]

OpenStudy (anonymous):

If so, then this is how I did it. 1) First, get that -7 over to the left side to make it a nice quadratic formula like so: \[2x ^{2} -8x +7 = 0\] 2) Now, if we try to factor this, it doesn't come out nice. So, if we can't flat out factor it, what's another method we can use to make a formula factor-able?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

(Hint: what can we COMPLETE to make this quadratic SQUARED formula factor-able?)

OpenStudy (anonymous):

I cant remember

OpenStudy (anonymous):

Complete the...

OpenStudy (anonymous):

square

OpenStudy (anonymous):

Ok, so, what's the formula for completing the square?

OpenStudy (anonymous):

i dont know my teacher doesn't explain anything.

OpenStudy (anonymous):

Ok, well, the formula for completing the square is: \[\frac{ b }{ 2a }\] with, \[ax ^{2} + bx + c = 0\] So, what's A and B?

OpenStudy (anonymous):

I should mention that A, B, and C are all constants, not just separate variables themselves.

OpenStudy (anonymous):

2 and 8

OpenStudy (anonymous):

Good, but we have an issue. The way this is currently worded has that pesky 2 out in front. That's no good! We need a formula like the following: \[a(x + d)^{2} + e = 0\] With \[d = \frac{ b }{ 2a }\] and \[e = c - \frac{ b ^{2} }{ 4a }\] So, what can we do to the formula to make \[2x ^{2} -8x+7=0\] look like: \[ax ^{2}+bx+c\] ?

OpenStudy (anonymous):

okay i got lost

OpenStudy (anonymous):

Im starting to think its the third one after looking at it for a long time

OpenStudy (anonymous):

I'll go ahead and continue, but try to work that out before looking ahead. To get it in proper format, we can divide everything by two, which leads us to: \[x ^{2}-4x+\frac{ 7 }{ 2 }=0\] Now we do completing the square, which leads to: \[(x-2)^{2} - \frac{ 1 }{ 2 } = 0\] Now, just some simple simplification. 1) \[(x-2)^{2} = -\frac{ 1 }{ 2 }\] 2) \[(x-2)=\sqrt{\frac{ 1 }{ 2 }}\] or \[(x-2)=-\sqrt{\frac{ 1 }{ 2 }}\] 3) \[x=2-\sqrt{\frac{ 1 }{ 2 }}\] or \[x=2+\sqrt{\frac{ 1 }{ 2 }}\]

OpenStudy (anonymous):

Oops! Had an arithmatic error in there. Before 1), I meant to move the - 1/2 to the right side.

OpenStudy (anonymous):

Did that make sense? If not, say so, and I'll find another way to explain it.

OpenStudy (anonymous):

i understand it now once i see the numbers

OpenStudy (anonymous):

Sorry, I know that seemed a bit confusing, but I promise, it's not as bad as it seems.